# Comparing Hallucination Detection Metrics for Multilingual Generation

### v1

@@ 4.2 Correlation Study Across Metrics
In this section, we conduct a correlation analysis to determine how various metrics align in measuring hallucination in multilingual contexts. This includes (1) the correlation between lexical hallucination metrics and NLI-based metrics, (2) the performance of the four reference-based NLI metrics, and (3) the relationship between pairwise metrics and reference-based metrics.
@@
Alessandro Del Piero
Gen: Alessandro Del Piero, born on September 28, 1976 in Brescia, Italy, is a former Italian professional football player who served as a forward. Wiki: Alessandro Del Piero, Italian male football player…The old Maldini, who was the head coach of the national team at the time, appointed newcomers Del Piero and Vieri as the main forwards… Comment: There are 4 facts in this sentence, with 1 contradictory hallucination and 1 unverifiable hallucination. The birth date is wrong; Wikipedia doesn’t mention the birth place; for the last entity, the evidence indirectly support it.
Example of Instruction-Conflict Hallucination
Carl Edward Sagan
Gen: Stephen Hawking (January 8, 1942 -) is a famous British physicist and cosmologist.
Example of Preceding-Context-Conflict Hallucination
Blanca of Castile
Gen: Blanca of Castile (Blanca of Castile), born in 1188 and died in 1252, was the princess of Castile, queen and regent of France. Blanca is a city in the Autonomous Community of Castilla-La Mancha, Spain.
Table 4: Examples of annotation process and types of hallucinations. Sentences highlighted in red represent verifiable hallucinations that contradict evidence found in Wikipedia; those marked in yellow denote unverifiable hallucinations which lack clear evidence for verification and content highlighted in green is considered factual.
@@
Figure 1 shows that in high-resource languages (i.e., English, Chinese, French, Spanish, Vietnamese, and Indonesian), ROUGE-1 and ROUGE-L metrics demonstrate a high degree of correlation, as do the ROUGE metrics and Named Entity Overlap (NEO). However, we generally find no correlation between lexical- and NLI-based metrics, indicating that while both lexical- and NLI-based approaches are commonly proposed as automatic methods for hallucination detection, they do not measure the same deviations from a reference text.
@@
We also observe interesting trends regarding the relationship between different NLI-based metrics (Figure 1). We find that $\mathrm{ENT}$ scores are highly correlated with the $\mathrm{DIFF}$ score, indicating that these metrics identify similar artifacts in the text. Surprisingly, no correlation is observed between $\mathrm{UNV}$ and $\mathrm{CON}$ scores. Moreover, we find a negative correlation between $\mathrm{UNV}$ and $\mathrm{CON}$ scores. This is because sentences that include verifiable hallucinations likely contradict the reference text. In contrast, sentences with information that is unsubstantiated by the reference (e.g., unverifiable) will be identified as neutral instead.
Pairwise and reference-based metrics correlate in higher-resource languages.
@@
For high-resource languages in the XNLI finetuning dataset (English, French, Chinese, Spanish, Bulgarian, and German), we observe a higher correlation in the range of 0.35 to 0.56 for pairwise NLI metrics when it comes to detecting verifiable hallucinations by $\mathrm{ENT}$ score (Table 2). This suggests pairwise metrics can identify generated content that deviates from the reference. However, the Pearson Correlation Coefficient shows lower correlation values (in the range of 0.15 to 0.21) compared to pairwise $\mathrm{UNV}$. For lower-resource languages, such as Finnish, Italian, Persian, and others, the correlation coefficients with the entailment score are often in the range of 0.00 to 0.30 and not statistically significant, highlighting the challenge of effective hallucination detection in contexts with limited language resources.
5 Human Evaluation of Multilingual Hallucination Metrics
Currently, there are no standard multilingual hallucination detection datasets available on which to compare the considered methods. We, therefore, annotate model generations in English and Chinese; annotations are performed by native speakers of each language. They manually find all verifiable and unverifiable hallucinations by checking if the generation is supported by the Wikipedia reference at both the sentence- and atomic-fact-level. Examples of resulting annotations are shown in Table 4.
l—cccc—cccc

### v2

@@ 4.2 Correlation Study Across Metrics
In this section, we conduct a correlation analysis to determine whether the considered metrics agree in measuring hallucination in multilingual contexts. This includes (1) the correlation between lexical hallucination metrics and NLI-based metrics, (2) the agreement of the four reference-based NLI metrics, and (3) the relationship between pairwise metrics and reference-based metrics.
@@
Alessandro Del Piero Gen: Alessandro Del Piero, born on September 28, 1976 in Brescia, Italy, is a former Italian professional football player who served as a forward. Wiki: Alessandro Del Piero, Italian male football player…The old Maldini, who was the head coach of the national team at the time, appointed newcomers Del Piero and Vieri as the main forwards… Comment: There are 4 facts in this sentence, with 1 contradictory hallucination and 1 unverifiable hallucination. The birth date is wrong; Wikipedia doesn’t mention the birth place; for the last entity, the evidence indirectly support it.
Table 4: Example of human hallucination annotations. Red represents verifiable hallucinations contradicting evidence in the reference (Wiki), yellow denotes unverifiable hallucinations without relevant evidence, and green is supported by the reference text.
@@
Figure 1 shows that in high-resource languages (i.e., English, Chinese, French, Spanish, Vietnamese, and Indonesian), ROUGE-1 and ROUGE-L metrics demonstrate a high degree of correlation, and Named Entity Overlap (NEO) correlates reasonably well with ROGUE precision metrics. However, we generally find no correlation between lexical- and NLI-based metrics, indicating that while both lexical- and NLI-based approaches are commonly proposed as automatic methods for hallucination detection, they do not measure the same deviations from a reference text.
@@
We also observe interesting trends regarding the relationship between different NLI-based metrics (bottom right-hand corner of Figure 1). We find that $\mathrm{ENT}$ scores are highly (inversely) correlated with the $\mathrm{DIFF}$ score, indicating that these metrics identify similar artifacts in the text. Moreover, we find a negative correlation between $\mathrm{UNV}$ and $\mathrm{CON}$ scores. This is because sentences that include verifiable hallucinations likely contradict the reference text. In contrast, sentences with information that is unsubstantiated by the reference (e.g., unverifiable) will be identified as neutral instead.
Pairwise and reference metrics do not correlate in low-resource languages.
@@
For high-resource languages in the XNLI finetuning dataset (English, French, Chinese, Spanish, Bulgarian, and German), we observe higher correlations ranging between 0.35 to 0.56 for pairwise and reference-based NLI metrics when it comes to detecting verifiable hallucinations by $\mathrm{ENT}$ score (Table 3). This suggests pairwise metrics can identify generated content that deviates from the reference and may be useful for detecting hallucinations when gold reference texts are not available. However, the Pearson Correlation Coefficient shows lower correlation values (in the range of 0.15 to 0.21) when comparing pairwise and reference-based $\mathrm{UNV}$, indicating a less effective capture of extrinsic hallucinations involving plausible yet unverifiable information. For lower-resource languages, such as Finnish, Italian, Persian, correlation with the entailment score is often not statistically significant. This implies that the effectiveness of pairwise hallucination metrics is limited to higher-resource languages, highlighting the challenge of effective hallucination detection in limited resource contexts.
Question Instructions
0. Atomic-Level Annotation Extract a sentence that contains only one simple fact.
1. Evidence Extraction Copy and paste all relevant evidence.
2. Total Facts Provide an approximate count of the total facts. Each date is counted as one fact, except for birthdates and death dates, which are counted as two separate facts.
3. Verifiable Correct Facts Count the number of facts that can be verified as correct.
4. Verifiable Contradictory Facts Count the number of facts that contradict verified information.
5. Unverifiable Facts Count the number of facts that cannot be verified.
6. Conflict with Preceding Context Indicate whether there is any conflict with the preceding context (True or False).
7. Conflict with Instructions Determine whether there is any conflict with the instructions. Label it as False if the example provides (a) a biography of (b) the correct person. Otherwise, label it as True.
Table 5: Instructions for manual hallucination annotations. Step 0 is only taken for atomic-level fact verification.
5 Human Evaluation
We manually annotate the model generations analyzed in the prior section; the annotations are performed on paired subsets of the English and Chinese generations by native speakers. Following the Attributable to Identified Sources (AIS) paradigm Rashkin et al. (2023) for measuring hallucination, annotators manually find all verifiable and unverifiable hallucinations by checking if the generated output is attributable to the Wikipedia reference at both the sentence- and atomic-fact-level. Table 4 shows example annotations.
{NiceTabular}
l|ccc|ccc
Sentence Level Atomic-Fact Level
[…]
Random - 10.84 82.86 - 52.1143.56
[…]
R1-P. 0.08  19.78 81.48 0.10 51.23 44.23
RL-P. 0.11  20.04 83.10 0.12  52.18 42.83
NEO-P. 0.14  17.49 80.84 0.09 53.09 45.32
$\mathrm{DIFF}$ 0.21 38.46 89.49 0.19 57.46 54.41
$\mathrm{ENT}$ 0.31 40.32 90.86 0.23 60.71 57.48
$\mathrm{CON}$ 0.11 16.47 80.41 -0.01 51.49 52.16
Reference
R1-P. 0.21 30.05 89.08 0.19 53.28 46.25
RL-P. 0.17 28.54 85.35 0.13  50.31 49.93
NEO-P. 0.17  16.15 83.75 0.12 57.54 47.51
$\mathrm{DIFF}$ 0.34 56.11 94.14 0.31 65.85 60.90
$\mathrm{ENT}$ 0.49 65.32 94.96 0.35 68.00 63.69
$\mathrm{CON}$ 0.08 31.56 87.49 -0.19 53.18 57.43
$\mathrm{mFact}$ 0.20 35.68 91.16 0.29 67.30 61.67
$\mathrm{Seahorse}$0.17  13.25 75.40 -0.07  53.30 46.67
Table 6: Comparison of sentence and atomic-fact verifiable hallucination metrics with the human support rate. F denotes factual examples and NF denotes non-factual examples. p-values of this correlation is larger than 0.05.
5.1 Experimental Setup
The authors (one per language) manually checked every sentence in the audited subset of generations, using the steps listed in Table 5. For atomic-fact-level annotation, a preprocessing step is taken to extract only sentences that contain a standalone proposition [^5] A standalone proposition is independently interpretable from the information contained in the assertion. Rashkin et al. (2023). Then, for both the sentence- and atomic-fact-level annotation, we annotate all relevant evidence sentences from the reference Wikipedia page and accumulate the counts for different types of propositions (Table 5). Table 9 details data statistics from this annotation process.
Metrics
We compare our automatic metrics presented in §2.2 with human annotations using correlation and classification; we specifically compare precision metrics because they are generally the strongest automatic measure in Section 4.1. For correlation, we investigate the relationship of the metrics with the support rate (SR; $N_{vs}/N_{t}$) for verifiable hallucination detection and with the unverified rate ($N_{nv}/N_{t}$) for unverifiable hallucination detection using their Pearson correlations.
To consider the classification agreement of these metrics, we calculate the Precision-Recall area under the curve (AUC-PR) between the discretized human-annotated and automatic metrics. We convert human annotations into classification labels by labeling an example as factual only if all its facts are supported by evidence for verifiable hallucinations with the support rate (Table 5); for unverifiable hallucinations, we consider any sentence with at least one fact not supported or refuted by the reference to be unverified: $N_{nv}\geq 1$ (Table 5.3).
We discretize the automatic NLI-based metrics into classification labels by setting their respective thresholds, with 0.5 for the entailment and contradictory scores and 0 for the difference between these two scores. The thresholds were selected based on the different degrees of tolerance for the proportions of unverifiable hallucinations in a sentence. We then perform classification using the discretized human judgments as gold labels.
5.2 Automatic Metric Results
NLI entailment outperforms lexical metrics on sentence-level verification.
We observe low correlation between lexical metrics like ROUGE-1 (R1) and Named Entity Overlap (NEO) and the human-annotated support rate (SR; Table 5). Moreover, AUC-PR for these metrics exhibits minor improvements over the random classification baseline, particularly for non-factual examples (NF), underscoring the limit of lexical metrics for accurately detecting factual hallucinations.
In contrast, the NLI-based metrics highly correlate with different measures of human verification for detecting verifiable hallucinations at the sentence level (Table 5). These results corroborate the NLI metric results in English in Manakul et al. (2023). Specifically, the $\mathrm{ENT}$ and $\mathrm{DIFF}$ scores perform well, while $\mathrm{CON}$ alone does not demonstrate the same agreement with human judgments.
Pairwise metrics underperform reference-based metrics.
When compared against human annotations, pairwise metrics always underperform the reference-based ones: their Pearson correlations, AUCF, AUCNF, and accuracies are all worse than the metrics using references, and this holds at both the sentence- and atomic-fact-levels (Table 5). However, we note that pairwise NLI-based metrics, particularly the entailment and difference scores, demonstrate significantly better performance than the lexical-based pairwise metrics, though they still underperform comparable reference metrics. This suggests pairwise NLI-based approaches remain the most useful approach in hallucination detection settings where references are unavailable.
NLI metrics struggle to detect unverifiable hallucination and check atomic facts.
Comparison against human judgments on unverifiable hallucination detection reveals that automatic metrics show only marginal improvements over random classification and low correlation with human annotations (Table 5.3). Furthermore, NLI metrics also encounter significant challenges in accurately verifying the factuality of simple atomic facts (Table 5). Both AUC-PR and accuracy demonstrate a marked decrease in the effectiveness of NLI-based metrics on atomic facts. This aligns with Luo et al. (2022), which also highlighted similar limitations in English NLI metrics for verification. This finding underscores the need for alternative metrics to address current NLI approaches’ limitations, especially in the critical area of atomic factuality evaluation.
5.3 Supervised Metric Results
We also consider two supervised metrics: mFACT Qiu et al. (2023) and seahorse-Q4 Clark et al. (2023). mFACT Qiu et al. (2023) automatically apply English hallucination detection approaches to new languages by first using faithfulness metrics to rank and annotate English examples, and then translating the most and least faithful samples into a set of target languages. They then train a classifier on each target language (including English and Chinese). The seahorse-Q4 metric Clark et al. (2023) similarly fine-tunes a mT5-large model Xue et al. (2021) on an attribution task, using attribution subset (Q4) of the SEAHORSE dataset.
Supervised metrics underperform NLI-based metrics in verifiable hallucination detection.
mFACT scores exhibit similar but slightly lower performance than NLI-based metrics at both the sentence and atomic fact levels (Table 5). The seahorse-Q4 approach shows comparable performance to mFACT in English (Appendix Table D). However, as seahorse-Q4 does not see Chinese data in training, it fails to detect verifiable hallucinations in this setting, leading to a lower overall score and highlighting the brittleness of supervised methods for hallucination detection.
Supervised metrics detect unverifiable hallucinations more effectively.
[…]
This trend also holds at the atomic-fact level, where the supervised metrics outperform other methods in across all measures. However, they still exhibit lower correlation with this unverified rate than at the sentence level, indicating that detecting hallucination on simpler atomic facts remains challenging in the unverified setting.
{NiceTabular}
[…]
$\mathrm{UNV}$${}_{\textit{Sent}}$ 0.28 31.77 71.84
$\mathrm{mFACT}$${}_{\textit{Sent}}$0.29 54.17 69.67
$\mathrm{seahorse}$${}_{\textit{Sent}}$0.10 42.08 69.49
[…]
$\mathrm{UNV}$${}_{\textit{Atom}}$ 0.12 77.49 27.36
$\mathrm{mFACT}$${}_{\textit{Atom}}$0.13 80.35 31.49
$\mathrm{seahorse}$${}_{\textit{Atom}}$0.09 78.32 28.48
Table 7: Agreement with human judgments on unverifiable hallucination detection, averaged for Chinese and English. UNV denotes unverifiable hallucination, and V denotes verifiable. Gray values have p-values $>$ 0.05.
6 Related Work
Factuality Hallucination Detection
Detecting hallucinations is crucial for ensuring the reliability of machine-generated content. One line of work leverages uncertainty by analyzing the probability of the LLM’s output space Mielke et al. (2022); Kadavath et al. (2022); Varshney et al. (2023); other methods evaluate consistency between repeatly generated samples (including NLI-based approaches) Elaraby et al. (2023); Manakul et al. (2023), similar to our pairwise evaluation setting. Similarly, our reference-based evaluations follow works that use external evidence as reference texts and verify if the generation is supported by this reference Chern et al. (2023); Min et al. (2023). However, factuality hallucination evaluations are limited in multilingual settings, with existing works using prompting methods Ahuja et al. (2023) and machine translation Lai et al. (2023).
Task-specific Multilingual Hallucination
The main focus of multilingual hallucination evaluation methods so far has been on specific downstream tasks. mFACT evaluates faithfulness in summaries by transferring English judgments into target languages via machine translation Qiu et al. (2023). Similarly, Aharoni et al. (2022) leverages factual consistency models to improve faithfulness in multilingual summarization. In neural machine translation (NMT), Dale et al. (2022) detect and alleviate hallucinations by measuring generation similarity to the source text; Xu et al. (2023) similarly study source effects on hallucination with input perturbations. Other works Lee et al. (2019); Raunak et al. (2021) analyze the susceptibility of current NMT methods to generate hallucinations.
7 Conclusion
This study investigates the effectiveness of automatic metrics for detecting factual hallucinations in non-English generations by considering these metrics directly (Section 4) and comparing their predictions to human judgments and supervised detection approaches (Section 5). We document that while traditional lexical metrics struggle to detect hallucinations in multilingual settings, NLI-based metrics show promise in high-resource languages at the sentence level. However, their effectiveness diminishes when applied to atomic facts, and the reliability of NLI-based metrics is tied to the performance of NLI models, posing a significant hurdle in lower-resource languages. Therefore, our analysis highlights that detecting hallucinations effectively in a language is directly linked to the availability and quality of linguistic resources in that language. As a result, automatically detecting hallucinations in lower-resource languages remains a significant challenge for current NLP methods.
Limitations
This study focuses on text generation and hallucination in a specific setting (namely, generating biographical summaries with the BLOOM-mt model) to perform a controlled study on how different automatic metrics detect factual hallucinations. It remains an open question whether these findings hold in other generation settings, particularly when there is less reliance on factual knowledge (such as story generation).
Additionally, portions of our experimental setup rely on automatic methods. Specifically, we use machine translation to construct the prompt templates, which may introduce noise. Furthermore, due to the unavailability of native speakers for other languages, our human evaluation and comparison against automated metrics is limited to Chinese and English. In the future, we would like to expand on these findings with human evaluations in other, lower-resourced languages to confirm how well the automatic detection methods hold up in these settings.
Appendix A Qualitative Analysis of Challenging Cases in Annotation
We identify four challenging categories of hallucination detection for annotators and NLI metrics (Table 8).
  • Inferred: Implicit fact connections between generation and evidence.
  • Subjective: Generation contains subjective content, which is challenging for both human annotators and fact-based NLI models.
  • Nuanced Difference: There are subtle distinctions between evidence and generated text, which is often missed by surface-level text classification in NLI models.
  • Temporal Information: Generation contains time-sensitive information, which requires models to have an understanding of temporal context.
Type Example
Inferred Gen: Alessandro Del Piero is a former Italian professional football player and a forward. Wiki: Maldini, who was the head coach of the national team, appointed Del Piero as the main forward.
Subjective Gen: Frida Kahlo is widely regarded as the most influential painter of the 20th century.
Nuanced Difference Gen: Louis Pasteur is known as the "Father of Modern Microbiology". Wiki: … has been honored as the "father of microbiology"…
Temporal Information Gen: Michelle Bachelet is currently the President of Chile. Wiki: She served as President of Chile from 2006 to 2010 and from 2014 to 2018…
Table 8: Categories of special case in annotation.
Each category presents unique difficulties in determining the factuality of generated content with its evidence source.
  • Inferred connection between generated content and evidence is one of the biggest challenges for both annotators and NLI models, since they need to infer the relationship or the factual basis that links them. This requires a deep understanding of context and the ability to draw inferences from potentially sparse or indirect evidence.
  • Subjective content in generations poses a significant challenge because it introduces personal opinions, emotions, or interpretations that are inherently difficult to verify against factual evidence. For human annotators, this can lead to variability in judgments based on personal biases or interpretations. For NLI models, which are primarily designed for fact-based analysis, handling subjective content requires advanced understanding of sentiment, opinion, and cultural context, areas where current models may fall short.
  • Nuanced difference between evidence and generated text highlight the limitations of surface-level text classification approaches in NLI models. Detecting nuanced differences demands a granular analysis of semantics, requiring models to understand context, synonyms, and slight variations in meaning. This challenge underscores the need for more sophisticated NLI models capable of deep semantic analysis and the importance of training annotators to pay attention to detail and understand the significance of minor discrepancies.
  • Time-sensitive information introduces complexity because it requires both annotators and models to have an understanding of temporal context and the ability to evaluate statements within the correct time frame. This can be particularly challenging when information changes over time, requiring up-to-date knowledge and the ability to discern the relevance of temporal qualifiers in text. For NLI models, this underscores the need for dynamic knowledge bases and the ability to reason about time, which are areas where current models may lack proficiency.
Overall, these challenges highlight the complexities involved in hallucination detection and the need for advanced capabilities in human annotators and NLI models.
Appendix B Metric Calculation Details
Appendix C Additional Results
We show the pairwise metric results in Table 10. We observe similar trends as the reference-based metrics. Also, the average statistics of annotation result are shown in Table 9.
Metric Sent-Level Atomic-Level
# Examples 111 102
# Words 46.21 10.21
# Evidence 2.17 1.00
# Total Facts 4.76 1.00
Support Rate 0.35 0.29
Contradictory Rate 0.15 0.24
Unverified Rate 0.50 0.47
Instruction-conflict Rate 0.03 0.07
Context-conflict Rate 0.13 0.06
Table 9: Average statistics of Chinese and English annotation data.
Table 10: Results of different pairwise consistency metrics for the BLOOMZ-mt model. "-" indicates they have no coverage for the NER tool we use. All of the ROUGE and Named Entity Overlap results are in percentage (%).
Language R1_F1 R1_P R1_R RL_F1 RL_P RL_R NEO_F1 NEO_P NEO_R $\mathrm{DIFF}$ $\mathrm{UNV}$
English 12.03 14.21 14.71 5.91 7.10 7.37 4.27 53.41 2.26 -0.25 0.57
Chinese 7.57 8.56 8.55 4.00 4.55 4.55 4.69 35.28 2.79 -0.29 0.57
Spanish 12.49 14.45 15.05 6.21 7.31 7.68 3.28 48.48 1.76 -0.22 0.52
German 4.42 10.47 4.48 1.98 5.51 2.04 0.83 36.06 0.42 -0.33 0.56
Russian 0.15 0.19 0.13 0.00 0.00 0.00 0.48 11.28 0.25 -0.09 0.55
Indonesian 4.74 6.23 6.52 1.59 2.13 2.27 - - - -0.16 0.68
Vietnamese 11.60 14.36 15.91 6.26 7.91 8.91 - - - -0.32 0.56
Persian 0.00 0.00 0.00 0.00 0.00 0.00 - - - -0.31 0.52
Ukrainian 0.00 0.00 0.00 0.00 0.00 0.00 0.70 20.64 0.36 0.14 0.42
Swedish 10.04 14.69 10.82 8.53 13.23 9.74 1.28 45.24 0.66 -0.20 0.54
Thai 0.01 0.02 0.01 0.00 0.00 0.00 0.00 0.00 0.00 -0.19 0.68
Japanese 0.47 0.75 0.58 0.08 0.13 0.09 0.47 15.52 0.25 -0.07 0.56
Romanian 12.21 13.95 12.68 8.57 9.40 8.70 0.39 17.47 0.20 -0.30 0.44
Hungarian 0.18 0.25 0.46 0.00 0.00 0.00 - - - -0.18 0.58
Bulgarian 0.30 0.44 0.28 0.05 0.08 0.06 - - - -0.02 0.96
French 12.02 14.03 14.61 6.26 7.43 7.76 4.35 57.41 2.31 -0.05 0.92
Finnish 0.46 0.47 0.61 0.17 0.19 0.19 0.58 23.71 0.30 -0.04 0.94
Korean 0.19 0.26 0.17 0.00 0.00 0.00 0.24 8.48 0.12 0.25 0.63
Italian 4.79 7.85 5.08 2.71 4.15 2.77 1.00 30.26 0.52 -0.01 0.97
{NiceTabular}
[…]
$\mathrm{UNV}$${}_{\textit{Sent}}$ 0.27 32.70 70.31
$\mathrm{mFACT}$${}_{\textit{Sent}}$
-0.18 45.2763.94
$\mathrm{seahorse}$${}_{\textit{Sent}}$-0.03 41.74 62.73
[…]
$\mathrm{UNV}$${}_{\textit{Atom}}$ 0.09 73.49 20.19
$\mathrm{mFACT}$${}_{\textit{Atom}}$0.17 79.03 32.01
$\mathrm{seahorse}$${}_{\textit{Atom}}$0.12 78.24 30.93
Table 11: Results of unverifiable hallucination detection in human evaluation in English, with the best agreement with humans indicated in bold. Sent denotes sentence-level detection, and Atom denotes atomic-fact-level detection. UNV denotes unverifiable hallucination, and V denotes verifiable. The number in gray have the p-values larger than 0.05.
{NiceTabular}
[…]
$\mathrm{UNV}$${}_{\textit{Sent}}$ 0.31 30.27 76.12
$\mathrm{mFACT}$${}_{\textit{Sent}}$0.32 56.54 60.42
$\mathrm{seahorse}$${}_{\textit{Sent}}$0.11 41.45 63.54
[…]
$\mathrm{UNV}$${}_{\textit{Atom}}$ 0.12 79.26 27.36
$\mathrm{mFACT}$${}_{\textit{Atom}}$0.19 86.77 34.38
$\mathrm{seahorse}$${}_{\textit{Atom}}$0.14 80.76 30.92
Table 12: Results of unverifiable hallucination detection in human evaluation in Chinese, with the best agreement with humans indicated in bold. All the p-values of the correlation is less than 0.05. Sent denotes sentence-level detection, and Atom denotes atomic-fact-level detection. UNV denotes unverifiable hallucination, and V denotes verifiable. The number in gray have the p-values larger than 0.05.
Appendix D Generation Prompt Templates
We present the full set of prompt templates for all languages from Section 3 in Figure 2.
Figure 2: Prompt templates of all languages used in generating biography. {} represents human names.
{NiceTabular}
l|ccc|ccc